# Zero-Shot Multi-Object Shape Completion

### v1

@@ Abstract
We present a 3D shape completion method that recovers the complete geometry of multiple objects in complex scenes from a single RGB-D image. Despite notable advancements in single object 3D shape completion, high-quality reconstructions in highly cluttered real-world multi-object scenes remains a challenge. To address this issue, we propose OctMAE, an architecture that leverages an Octree U-Net and a latent 3D MAE to achieve high-quality and near real-time multi-object shape completion through both local and global geometric reasoning. Because a naive 3D MAE can be computationally intractable and memory intensive even in the latent space, we introduce a novel occlusion masking strategy and adopt 3D rotary embeddings, which significantly improves the runtime and shape completion quality. To generalize to a wide range of objects in diverse scenes, we create a large-scale photorealistic dataset, featuring a diverse set of $12$K 3D object models from the Objaverse dataset which are rendered in multi-object scenes with physics-based positioning. Our method outperforms the current state-of-the-art on both synthetic and real-world datasets and demonstrates a strong zero-shot capability.
Figure 1: Given an RGB-D image and the foreground mask of multiple objects not seen during training, our method predicts their complete 3D shapes quickly and accurately, including occluded areas. (Left) Synthetic image results. (Right) Zero-shot generalization to a real-world image of household objects with noisy depth data. Our 3D results are rotated with respect to the input to highlight completions in occluded regions.
@@ 1 Introduction
Prior works [72, 44, 48, 35, 32, 37] have achieved phenomenal progress in scene and object shape completion from a single RGB-D image. Object-centric methods [17, 26] in particular can achieve very high reconstruction accuracy by relying on category-specific shape priors. However, when deployed on entire scenes such methods require bespoke instance detection/segmentation models, and often perform test-time optimization which is time consuming and would hinder real-time deployment on a robot. Moreover, existing methods are typically limited to a small set of categories. Thus, zero-shot 3D shape completion remains a challenging and open problem that has seen little success to date. This is in stark contrast to the sudden increase in powerful algorithms for 2D computer vision tasks such as object detection [76, 34] and image segmentation [71, 36]. We attribute this progress to a great extent to the availability of large-scale datasets [55, 8] coupled with neural architectures and learning objectives [51, 54, 22, 58] that can effectively exploit the highly structured data occurring in the natural world [20].
Taking inspiration from the latest developments in the 2D domain, we propose a shape completion algorithm at the scene level that generalizes across a large number of shapes and that only supposes an RGB-D image and foreground mask as input. Our method consists of Octree masked autoencoders (OctMAE) — a hybrid architecture of Octree U-Net and a latent 3D MAE (Figure 2). Although a recent work, VoxFormer [35], also extends MAE architecture to 3D using deformable 3D attention and shows great improvement in semantic scene completion tasks, its memory utilization is still prohibitive to handle a higher resolution voxel grid. We address this issue by integrating 3D MAE into the latent space of Octree U-Net. Our experiments show that the latent 3D MAE is the key to global structure understanding and leads to strong performance and generalization across all datasets. Moreover, we find that a choice of a masking strategy and 3D positional enbeddings is crucial to achieve better performance. We provide extensive ablations to verify that our design of 3D latent MAE is effective.
Our second contribution consists of the creation of a novel synthetic dataset to counteract the lack of large-scale and diverse 3D datasets. The dataset contains $12$K 3D models of hand-held objects from Objaverse [12] and GSO [16] datasets (Figure 4). We utilize the dataset to conduct a comprehensive evaluation of our method as well as other baselines and show that our method scales and achieves better results. Finally, we perform zero-shot evaluations on synthetic as well as real datasets and show that a combination of 3D diversity coupled with an appropriate architecture is key to generalizable shape completion in the wild.
@@
  • We create the first large-scale and diverse synthetic dataset using Objaverse [12] dataset for zero-shot multi-object shape completion, and provide a wide range of benchmark and analysis.
@@ 2 Related Work
Reconstructing indoor scenes and objects from a noisy point cloud has been widely explored [43, 49, 24, 9, 2, 10, 4, 6, 66, 25, 35, 67, 1, 57, 41, 48]. Several works [73, 72, 44, 5, 61, 64, 59, 75, 45, 4, 77, 48] tackle more challenging shape completion tasks where large parts of a target is missing. While these methods achieve impressive results, they do not explicitly consider semantic information, which may limit their capability for accurate shape completion. Recent methods [32, 35, 77, 33] in Semantic Scene Completion (SSC) leverage semantic information via an RGB image. Nevertheless, the number of target categories is quite limited, restricting its utility for a broad range of applications in the real world. In addition, many methods adopt occupancy or SDF as an output representation, which necessitates post-processing such as the marching cubes [42] and sphere tracing to extract an explicit surface. As another direction, GeNVS[3], Zero-1-to-3 [40], and 3DiM [65] explore single-view 3D reconstruction via novel view synthesis. However, expensive test-time optimization is required. Recently, One-2-3-45 [39] and MCC [67] attempt to improve the generation speed, however, their runtime for multi-object scenes is still far from near real-time. Further, since these methods are object-centric, multiple objects in a single scene are not handled well due to the complicated geometric reasoning especially caused by occlusions by other objects. In this paper, we propose a general and near real-time framework for multi-object 3D shape completion in the wild using only an RGB-D image and foreground mask without expensive test-time optimization.
@@
Recently, various types of implicit 3D representation have become popular in 3D reconstruction and completion tasks. Early works [48, 43, 18] use a one-dimensional latent feature to represent a 3D shape as occupancy and SDF fields. Several works [49, 32, 59] employ voxels, ground-planes, and triplanes, demonstrating that the retention of geometric information using 3D CNNs enhances performance. Although the voxel representation typically performs well among these three, its cubic memory and computational costs make it challenging to increase its resolution. To mitigate this issue, sparse voxels [63, 6, 56, 21, 38] treat a 3D representation as a sparse set of structured points using the octree and hash table and perform convolutions only on non-empty voxels and its neighbors. Further, the high-resolution sparse voxel enables a direct prediction of a target surface. As another direction, [1, 78, 68] leverage point cloud. Nonetheless, an unstructured set of points can be non-uniformly distributed in the 3D space and requires running the k-NN algorithm at every operation. This aspect often renders point-based methods less appealing compared to the sparse voxel representation. Therefore, our method adopts an octree-based representation used in [63] for efficient training and direct surface prediction.
@@ 3 Proposed Method
Given an RGB image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, depth map $\mathbf{D}\in\mathbb{R}^{H\times W}$, and foreground mask $\mathbf{M}\in\mathbb{R}^{H\times W}$ containing all objects of interest, we aim to predict their complete 3D shapes quickly and accurately. Our framework first encodes an RGB image $\mathbf{I}$ with a pre-trained image encoder $E$ such as ResNext and then lifts the resulting features up to 3D space using a depth map $\mathbf{D}$ and foreground mask $\mathbf{M}$ to acquire 3D point cloud features $\mathbf{F}\in\mathbb{R}^{N\times D}$ and its locations $\mathbf{P}\in\mathbb{R}^{N\times 3}$ (Section 3.1). Second, we convert the 3D features into an octree using the same algorithm used in [64] and pass it to OctMAE to predict a surface at each LoD (Section 3.2). The diagram of our method is visualized in Figure 2.
@@ 3.2 OctMAE: Octree Masked Autoencoders
We design OctMAE which leverages Octree U-Net [63] and latent 3D MAE to achieve accurate and efficient shape completion. Octree U-Net consists of multiple sparse 3D convolutional layers. While they can efficiently encode octree features to low resolution, only local regions are considered at each operation. On the contrary, 3D MAE can capture global object information which helps predict globally consistent 3D shapes. However, unlike an image, a dense voxel grid contains a prohibitive number of tokens even in the latent space, which makes it challenging to adopt an MAE architecture directly for 3D tasks. Recently, ConvMAE [19] proposed to leverage the advantages of both CNNs and MAE in 2D for efficient training. Nevertheless, a naïve extension of ConvMAE [19] to 3D also leads to prohibitive computational and memory costs. To address this issue, we propose a novel occlusion masking strategy and adopt 3D rotary embeddings, enabling efficient masked autoencoding in the latent space.
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@ 4 Dataset
As shown in Table 1, existing datasets are limited in the diversity of object categories. Although the CO3D V2 dataset [53] contains data for 40k objects, because the provided ground-truth 3D shapes are reconstructed from unposed multi-view images, they tend to be highly noisy and parts of the object missing due to lack of visibility. To tackle this problem, we leverage Objaverse [12], a large-scale $1$M 3D object dataset containing $46$k objects with LVIS category annotations. To focus on shape completion of hand-held objects, we select $601$ categories and ensure that the largest dimension of the objects in each category approximately falls within the range of $4$ cm to $40$ cm. In addition, for high-quality rendering, we omit objects that lack textures, contain more than $10,000$ vertices, or are articulated. To increase the number of objects, we add objects from Google Scanned Objects (GSO) [16], which results in $12,655$ objects in total. We render $1$M images of $25,000$ scenes using physics-based rendering and positioning via BlenderProc [13] to simulate realistic scenes (Figure 4). For each image, we randomly choose a camera view such that at least one object is within the camera frame. We also generate $1,000$ images using $250$ withheld objects for evaluation.
@@ 5 Experimental Results
We evaluate the baselines and our model on one synthetic and three real-world datasets. For the synthetic dataset, we render $1,000$ images using textured 3D scans from Objaverse [12], following the same procedure described in Section 4. We randomly choose $3$ to $5$ objects per image from the withheld objects for Objavese dataset. Since these 3D scans are relatively more complex than the objects seen in the real-world datasets we use, they can provide a good shape completion quality estimate for complex objects. For the real-world dataset, we use the YCB-Video [69], HOPE [37] and HomebrewedDB (HB) [29] datasets. YCB-Video consists of $21$ everyday objects with diverse shapes. HOPE contains $28$ simple household objects with mostly rectangular and cylindrical everyday shapes, and the images are captured in various lighting conditions in indoor scenes using a RealSense D415 RGBD camera. HB includes $33$ objects (e.g., toy, household, and industrial objects). Their images are taken by PrimeSense Carmine in lab-like environments.
@@
As discussed in Secs. 1 and 2, multi-object shape completion from a single RGB-D image is relatively not explored due to the lack of large-scale and diverse multi-object shape completion datasets. We carefully choose baseline architectures that can support this task with simple or no adaptation. We focus on three primary method types from related fields. Firstly, we select Semantic Scene Completion (SSC) methods [32, 64, 6, 35] that do not heavily rely on domain or categorical knowledge of indoor or outdoor scenes. Secondly, we opt for object shape completion methods [6, 67, 64, 72] that can be extended to multi-object shape completion without an architectural modification and prohibitive memory utilization. Thirdly, we consider voxel or octree-based 3D reconstruction methods [49, 64, 6, 1] that predict a complete and plausible shape using noisy and sparse point cloud data. For dense voxel-based (e.g., AICNet [32], ConvONet [49] and VoxFormer [35]) and sparse voxel-based methods (e.g., MinkowskiNet [6], OCNN [64], and our method), we use LoD-$6$ and LoD-$9$ as an input resolution respectively. All the experiments are conducted using the original implementation provided by the authors, with few simple modifications to adapt for multi-object shape completion and a fair comparison. For instance, we extend the baselines that take the point cloud as input by concatenating the image features to the point cloud features. For occupancy-based methods, though their output voxel grid resolution is LoD-$6$, we use trilinear interpolation to predict occupancy at LoD-$7$ [49]. For MinkowskiNet [6] and OCNN [63, 64], we use the U-Net architecture with the depth of $5$ (LoD-$9$ to LoD-$4$). We discuss further details about the baseline architectures, their modifications, and hyperparameters in the supplemental.
Table 2: Quantitative evaluation of multi-object shape completion on Ours, YCB-Video [69], HOPE [37], and HomebrewedDB [29] datasets. Chamfer distance (CD), F1-Score@10mm (F1), and normal consistency (NC) are reported. Chamfer distance is reported in the unit of mm.
@@
Minkowski [6] Sparse 11.47 0.746 0.802 8.04 0.761 0.717 8.81 0.728 0.719 8.18 0.731 0.706
OCNN [64] Sparse 9.05 0.782 0.828 7.10 0.778 0.771 7.02 0.792 0.736 7.92 0.740 0.730
Ours Sparse 6.71 0.831 0.840 6.34 0.797 0.773 6.18 0.819 0.760 6.40 0.799 0.737
@@ 5.1 Quantitative Results
Table 2 shows that our method outperforms the baselines on all the metrics and datasets. Although our model is only trained on synthetic data, it demonstrates strong generalizability to real-world datasets. We also remark that our method exhibits robustness to the noise characteristics present in depth data captured by typical RGB-D cameras despite being trained on noise-free depth data in simulation. The comparisons show that hierarchical structures and the latent 3D MAE are key to predicting 3D shapes of unseen objects more accurately than the baselines. Unlike our method, VoxFormer [35] uses an MAE with 3D deformable attention where only $8$ neighbors of the reference points at the finest resolution are considered. Figure 8 also demonstrates that methods using a dense voxel grid or implicit representation fail to generalize to novel shapes. This implies that capturing a right choice of a network architecture is crucial to learn generalizable shape priors for 3D shape completion. Our method has the similar U-Net architecture used in MinkowskiNet [6] and OCNN [63] except we use the latent 3D MAE at LoD-$5$ instead of making the network deeper. This indicates that the latent 3D MAE can better approximate the shape distribution of the training dataset by leveraging an attention mechanism to capture a wider range of 3D contexts. Table 7 also confirms that our method achieves the best shape completion quality by measuring Chamfer distance in visible and occluded regions separately.
@@
3D DSA [35] 12.14 0.709 93.3
Neighbor. Attn. [78] 8.48 0.731 130.8
Octree Attn. [62] 7.66 0.752 116.4
Neighbor. Attn. [78] 8.39 0.753 111.9
Octree Attn. [62] 7.25 0.770 105.3
Full + Self Attn. 6.63 0.786 86.2
Full + Cross Attn. 6.40 0.799 85.1
@@
As shown in Table 4, we explore the effect of RoPE [60] on the validation set of our synthetic dataset. The first row shows that all the metrics significantly drop if positional encoding is not used. In addition, we test CPE [7] and APE [35] and obtain slightly better scores. CPE [7] is typically more effective than APE in tasks such as 3D instance/semantic segmentation and object detection where a complete 3D point cloud is given. However, this result highlights the challenge of capturing position information from mask tokens which initially have the identical parameters. Our method employs RoPE [60] for relative positional embedding. One of the important aspect of RoPE [60] is that it does not have any learnable parameters. Despite this, it demonstrates superior performance compared to other approaches. Although RoPE is originally proposed in the domain of natural language processing, our experiment reveals its effectiveness in multi-object 3D shape completion.
@@
OctFormer [62] 7.44 0.753 0.731 114.4
Octree U-Net [63] 6.18 0.819 0.760 85.1
Table 7: Comparisons of the runtime (ms) for shape completion. For reference, we also show Chamfer distance of visible $\text{CD}_{vis}$ and occluded $\text{CD}_{occ}$ regions on our synthetic dataset.
@@
We further explore the design of 3D latent MAE in Table 6. Increasing the number of layers in 3D latent MAE improves the shape completion quality while making the runtime slower. Consequently, we select $3$ layers for a good trade-off between the accuracy and runtime.
@@
In Table 6, we investigate U-Net architectures. The key difference of Minkowski U-Net [6] is the use a sparse tensor as an underlying data structure instead of an octree, which gives a slightly lower score than Octree U-Net [63]. OctFormer [62] proposes an octree-based window attention mechanism using the 3D Z-order curve to support a much larger kernel size than Octree U-Net. In general, a wider range of an effective receptive field helps achieve better performance. Nonetheless, OctFormer achieves a chamfer distance and F-1 score of $7.44$ and $0.753$, which is worse than Octree U-Net by $1.26$ and $0.66$ respectively. This indicates that the OctFormer’s attention mechanism is less effective compared to an Octree U-Net architecture especially in the presence of latent 3D MAE, playing the similar role in the latent space.
@@
Table 7 shows the runtime performance of the baselines and our method. For a fair comparison, we run inference over the $50$ samples of the HOPE dataset and report the average time. For occupancy-based methods, we predict occupancy on object surfaces and occluded regions. Due to the memory-intensive nature of MCC [1]’s Transformer architecture, we run inference multiple times with the maximum chunk size of $10,000$ points. Our experiments demonstrate that implicit 3D representations used in POCO [1] and MCC [67] become slower when the voxel grid resolution is higher. Further, an autoregressive Transformer adopted in ShapeFormer [72] greatly increases the runtime. Conversely, the methods which leverage sparse voxel grids (e.g., MinkowskiNet [6], OCNN [64], and Ours) achieve much faster runtime thanks to efficient sparse 3D convolutions, and hierarchical pruning on predicted surfaces. Our method offers runtimes comparable to the fastest method, while implementing attention operations over the scene via latent 3D MAE, and achieving superior reconstruction quality.
@@
Figure 6: Qualitative comparison of OCNN [63] and our method. Our proposed latent 3D MAE helps predict globally consistent shape completion.

### v2

# Zero-Shot Multi-Object Scene Completion

@@ Abstract
We present a 3D scene completion method that recovers the complete geometry of multiple unseen objects in complex scenes from a single RGB-D image. Despite notable advancements in single-object 3D shape completion, high-quality reconstructions in highly cluttered real-world multi-object scenes remains a challenge. To address this issue, we propose OctMAE, an architecture that leverages an Octree U-Net and a latent 3D MAE to achieve high-quality and near real-time multi-object scene completion through both local and global geometric reasoning. Because a naive 3D MAE can be computationally intractable and memory intensive even in the latent space, we introduce a novel occlusion masking strategy and adopt 3D rotary embeddings, which significantly improve the runtime and scene completion quality. To generalize to a wide range of objects in diverse scenes, we create a large-scale photorealistic dataset, featuring a diverse set of $12$K 3D object models from the Objaverse dataset that are rendered in multi-object scenes with physics-based positioning. Our method outperforms the current state-of-the-art on both synthetic and real-world datasets and demonstrates a strong zero-shot capability. https://sh8.io/#/oct_mae
Figure 1: Given an RGB-D image and the foreground mask of multiple objects not seen during training, our method predicts their complete 4D shapes quickly and accurately, including occluded areas. (Left) Synthetic image results. (Right) Zero-shot generalization to a real-world image of household objects with noisy depth data. Our 3D results are rotated with respect to the input to highlight completions in occluded regions.
@@ 1 Introduction
Prior works [74, 45, 49, 36, 33, 38] have achieved phenomenal progress in scene and object shape completion from a single RGB-D image. Object-centric methods [17, 26] in particular can achieve very high reconstruction accuracy by relying on category-specific shape priors. However, when deployed on entire scenes such methods require bespoke instance detection/segmentation models, and often perform test-time optimization which is time consuming and would hinder real-time deployment on a robot. Moreover, existing methods are typically limited to a small set of categories. Thus, zero-shot multi-object scene completion remains a challenging and open problem that has seen little success to date. This is in stark contrast to the sudden increase in powerful algorithms for 2D computer vision tasks such as object detection [78, 35] and image segmentation [73, 37]. We attribute this progress to a great extent to the availability of large-scale datasets [57, 8] coupled with neural architectures and learning objectives [52, 56, 22, 60] that can effectively exploit the highly structured data occurring in the natural world [20].
Taking inspiration from the latest developments in the 2D domain, we propose a scene completion algorithm at the scene level that generalizes across a large number of shapes and that only supposes an RGB-D image and foreground mask as input. Our method consists of Octree masked autoencoders (OctMAE) — a hybrid architecture of Octree U-Net and a latent 3D MAE (Figure 2). Although a recent work, VoxFormer [36], also extends MAE architecture to 3D using deformable 3D attention and shows great improvement in semantic scene completion tasks, its memory utilization is still prohibitive to handle a higher resolution voxel grid. We address this issue by integrating 3D MAE into the latent space of Octree U-Net. Our experiments show that the latent 3D MAE is the key to global structure understanding and leads to strong performance and generalization across all datasets. Moreover, we find that the choice of a masking strategy and 3D positional embeddings is crucial to achieve better performance. We provide extensive ablations to verify that our 3D latent MAE design is effective.
Our second contribution consists of the creation of a novel synthetic dataset to counteract the lack of large-scale and diverse 3D datasets. The dataset contains $12$K 3D models of hand-held objects from Objaverse [12] and GSO [16] datasets (Figure 4). We utilize the dataset to conduct a comprehensive evaluation of our method as well as other baselines and show that our method scales and achieves better results. Finally, we perform zero-shot evaluations on synthetic as well as real datasets and show that a combination of 3D diversity coupled with an appropriate architecture is key to generalizable scene completion in the wild.
@@
  • We create the first large-scale and diverse synthetic dataset using Objaverse [12] dataset for zero-shot multi-object scene completion, and provide a wide range of benchmark and analysis.
@@ 2 Related Work
Reconstructing indoor scenes and objects from a noisy point cloud has been widely explored [44, 50, 23, 9, 2, 10, 4, 6, 68, 24, 36, 69, 1, 59, 42, 49]. Several works [75, 74, 45, 5, 63, 66, 61, 77, 46, 4, 79, 49] tackle more challenging shape completion tasks where large parts of a target is missing. While these methods achieve impressive results, they do not explicitly consider semantic information, which may limit their capability for accurate shape completion. Recent methods [33, 36, 79, 34] in Semantic Scene Completion (SSC) leverage semantic information via an RGB image. Nevertheless, the number of target categories is quite limited, restricting its utility for a broad range of applications in the real world. In addition, many methods adopt occupancy or SDF as an output representation, which necessitates post-processing such as the marching cubes [43] and sphere tracing to extract an explicit surface. As another direction, GeNVS[3], Zero-1-to-3 [41], and 3DiM [67] explore single-view 3D reconstruction via novel view synthesis. However, expensive test-time optimization is required. Recently, One-2-3-45 [40] and MCC [69] attempt to improve the generation speed, however, their runtime for multi-object scenes is still far from near real-time. Further, since these methods are object-centric, multiple objects in a single scene are not handled well due to the complicated geometric reasoning especially caused by occlusions by other objects. In this paper, we propose a general and near real-time framework for multi-object 3D scene completion in the wild using only an RGB-D image and foreground mask without expensive test-time optimization.
@@
Recently, various types of implicit 3D representation have become popular in 3D reconstruction and completion tasks. Early works [49, 44, 18] use a one-dimensional latent feature to represent a 3D shape as occupancy and SDF fields. Several works [50, 33, 61] employ voxels, ground-planes, and triplanes, demonstrating that the retention of geometric information using 3D CNNs enhances performance. Although the voxel representation typically performs well among these three, its cubic memory and computational costs make increasing resolution challenging. To mitigate this issue, sparse voxels [65, 6, 58, 21, 39] treat a 3D representation as a sparse set of structured points using the octree and hash table and perform convolutions only on non-empty voxels and its neighbors. Further, the high-resolution sparse voxel enables a direct prediction of a target surface. As another direction, [1, 80, 70] leverage point cloud. Nonetheless, an unstructured set of points can be non-uniformly distributed in the 3D space and requires running the k-NN algorithm at every operation. This aspect often renders point-based methods less appealing compared to the sparse voxel representation. Therefore, our method adopts an octree-based representation used in [65] for efficient training and direct surface prediction.
@@ 3 Proposed Method
Given an RGB image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, depth map $\mathbf{D}\in\mathbb{R}^{H\times W}$, and foreground mask $\mathbf{M}\in\mathbb{R}^{H\times W}$ containing all objects of interest, we aim to predict their complete 3D shapes quickly and accurately. Our framework first encodes an RGB image $\mathbf{I}$ with a pre-trained image encoder $E$ such as ResNeXt [72] and then lifts the resulting features up to 3D space using a depth map $\mathbf{D}$ and foreground mask $\mathbf{M}$ to acquire 3D point cloud features $\mathbf{F}\in\mathbb{R}^{N\times D}$ and its locations $\mathbf{P}\in\mathbb{R}^{N\times 3}$ (Section 3.1). Second, we convert the 3D features into an octree using the same algorithm used in [66] and pass it to OctMAE to predict a surface at each LoD (Section 3.2). The diagram of our method is visualized in Figure 2.
@@ 3.2 OctMAE: Octree Masked Autoencoders
We design OctMAE which leverages Octree U-Net [65] and latent 3D MAE to achieve accurate and efficient zero-shot multi-object scene completion. Octree U-Net consists of multiple sparse 3D convolutional layers. While the Octree U-Net architecture can efficiently encode octree features to low resolution, only local regions are considered at each operation. On the contrary, 3D MAE can capture global object information which helps predict globally consistent 3D shapes. However, unlike an image, a dense voxel grid contains a prohibitive number of tokens even in the latent space, which makes it challenging to adopt an MAE architecture directly for 3D tasks. Recently, ConvMAE [19] proposed to leverage the advantages of both CNNs and MAE in 2D for efficient training. Nevertheless, a naïve extension of ConvMAE [19] to 3D also leads to prohibitive computational and memory costs. To address this issue, we propose a novel occlusion masking strategy and adopt 3D rotary embeddings, enabling efficient masked autoencoding in the latent space.
@@ 4 Dataset
As shown in Table 1, existing datasets are limited in the diversity of object categories. Although the CO3D V2 dataset [54] contains data for 40k objects, because the provided ground-truth 3D shapes are reconstructed from unposed multi-view images, they tend to be highly noisy and parts of the object missing due to lack of visibility. To tackle this problem, we leverage Objaverse [12], a large-scale $1$M 3D object dataset containing $46$k objects with LVIS category annotations. To focus on completion of hand-held objects, we select $601$ categories and ensure that the largest dimension of the objects in each category falls approximately within the range of $4$ cm to $40$ cm. In addition, for high-quality rendering, we omit objects that lack textures, contain more than $10,000$ vertices, or are articulated. To increase the number of objects, we add objects from Google Scanned Objects (GSO) [16], which results in $12,655$ objects in total. We render $1$M images of $25,000$ scenes using physics-based rendering and positioning via BlenderProc [13] to simulate realistic scenes (Figure 4). For each image, we randomly choose a camera view such that at least one object is within the camera frame. We also generate $1,000$ images using $250$ withheld objects for evaluation.
@@ 5 Experimental Results
We evaluate the baselines and our model on one synthetic and three real-world datasets. For the synthetic dataset, we render $1,000$ images using textured 3D scans from Objaverse [12], following the same procedure described in Section 4. We randomly choose $3$ to $5$ objects per image from the withheld objects for Objavese dataset. Since these 3D scans are relatively more complex than the objects seen in the real-world datasets we use, they can provide a good scene completion quality estimate for complex objects. For the real-world dataset, we use the YCB-Video [71], HOPE [38] and HomebrewedDB (HB) [29] datasets. YCB-Video consists of $21$ everyday objects with diverse shapes. HOPE contains $28$ simple household objects with mostly rectangular and cylindrical everyday shapes, and the images are captured in various lighting conditions in indoor scenes using a RealSense D415 RGBD camera. HB includes $33$ objects (e.g., toy, household, and industrial objects). Their images are taken by PrimeSense Carmine in lab-like environments.
@@
As discussed in Secs. 1 and 2, multi-object scene completion from a single RGB-D image is relatively not explored due to the lack of large-scale and diverse multi-object scene completion datasets. We carefully choose baseline architectures that can support this task with simple or no adaptation. We focus on three primary method types from related fields. Firstly, we select Semantic Scene Completion (SSC) methods [33, 66, 6, 36] that do not heavily rely on domain or categorical knowledge of indoor or outdoor scenes. Secondly, we opt for object shape completion methods [6, 69, 66, 74] that can be extended to multi-object scene completion without an architectural modification and prohibitive memory utilization. Thirdly, we consider voxel or octree-based 3D reconstruction methods [50, 66, 6, 1] that predict a complete and plausible shape using noisy and sparse point cloud data. For dense voxel-based (e.g., AICNet [33], ConvONet [50] and VoxFormer [36]) and sparse voxel-based methods (e.g., MinkowskiNet [6], OCNN [66], and our method), we use LoD-$6$ and LoD-$9$ as an input resolution respectively. All the experiments are conducted using the original implementation provided by the authors, with few simple modifications to adapt for multi-object scene completion and a fair comparison. For instance, we extend the baselines that take the point cloud as input by concatenating the image features to the point cloud features. For occupancy-based methods, though their output voxel grid resolution is LoD-$6$, we use trilinear interpolation to predict occupancy at LoD-$7$ [50]. For MinkowskiNet [6] and OCNN [65, 66], we use the U-Net architecture with the depth of $5$ (LoD-$9$ to LoD-$4$). We discuss further details about the baseline architectures, their modifications, and hyperparameters in the supplemental.
Table 2: Quantitative evaluation of multi-object scene completion on Ours, YCB-Video [71], HOPE [38], and HomebrewedDB [29] datasets. Chamfer distance (CD), F1-Score@10mm (F1), and normal consistency (NC) are reported. Chamfer distance is reported in the unit of mm.
@@
Minkowski [6] Sparse 11.47 0.746 0.802 8.04 0.761 0.717 8.81 0.728 0.719 8.56 0.734 0.709
OCNN [66] Sparse 9.05 0.782 0.828 7.10 0.778 0.771 7.02 0.792 0.736 8.05 0.742 0.736
Ours Sparse 6.48 0.839 0.848 6.40 0.800 0.785 6.14 0.819 0.770 6.97 0.803 0.750
@@ 5.1 Quantitative Results
Table 2 shows that our method outperforms the baselines on all the metrics and datasets. Although our model is only trained on synthetic data, it demonstrates strong generalizability to real-world datasets. We also remark that our method exhibits robustness to the noise characteristics present in depth data captured by typical RGB-D cameras despite being trained on noise-free depth data in simulation. The comparisons show that hierarchical structures and the latent 3D MAE are key to predicting 3D shapes of unseen objects more accurately than the baselines. Unlike our method, VoxFormer [36] uses an MAE with 3D deformable attention where only $8$ neighbors of the reference points at the finest resolution are considered. Figure 8 also demonstrates that methods using a dense voxel grid or implicit representation fail to generalize to novel shapes. This implies that capturing a right choice of a network architecture is crucial to learn generalizable shape priors for zero-shot multi-object scene completion. Our method has the similar U-Net architecture used in MinkowskiNet [6] and OCNN [65] except we use the latent 3D MAE at LoD-$5$ instead of making the network deeper. This indicates that the latent 3D MAE can better approximate the shape distribution of the training dataset by leveraging an attention mechanism to capture global 3D contexts. Table 7 also confirms that our method achieves the best scene completion quality by measuring Chamfer distance in visible and occluded regions separately.
@@
3D DSA [36] 12.14 0.703 93.3
Neighbor. Attn. [80] 9.26 0.727 130.8
Octree Attn. [64] 7.99 0.752 116.4
Neighbor. Attn. [80] 8.81 0.759 111.9
Octree Attn. [64] 7.54 0.772 105.3
Full + Self Attn. 7.21 0.785 86.2
Full + Cross Attn. 6.97 0.803 85.1
@@
As shown in Table 4, we explore the effect of RoPE [62] on the validation set of our synthetic dataset. The first row shows that all the metrics significantly drop if positional encoding is not used. In addition, we test CPE [7], APE [36], and RPE [64] and obtain slightly better scores. CPE [7] is typically more effective than APE in tasks such as 3D instance/semantic segmentation and object detection where a complete 3D point cloud is given. However, this result highlights the challenge of capturing position information from mask tokens which initially have the identical parameters. Our method employs RoPE [62] for relative positional embedding. One of the important aspect of RoPE [62] is that it does not have any learnable parameters. Despite this, it demonstrates superior performance compared to other approaches. Although RoPE was originally proposed in the domain of natural language processing, our experiment reveals its effectiveness in multi-object 3D scene completion.
@@
OctFormer [64] 7.45 0.756 0.728 114.4
Octree U-Net [65] 6.14 0.819 0.770 85.1
Table 7: Comparisons of the runtime (ms). For reference, we also show Chamfer distance of visible $\text{CD}_{vis}$ and occluded $\text{CD}_{occ}$ regions on our synthetic dataset.
@@
We further explore the design of 3D latent MAE in Table 6. Increasing the number of layers in 3D latent MAE improves the scene completion quality while making the runtime slower. Consequently, we select $3$ layers for a good trade-off between the accuracy and runtime.
@@
In Table 6, we investigate U-Net architectures. The key difference of Minkowski U-Net [6] is the use of a sparse tensor as an underlying data structure instead of an octree, which gives a slightly better performance than Octree U-Net [65]. OctFormer [64] proposes an octree-based window attention mechanism using the 3D Z-order curve to support a much larger kernel size than Octree U-Net. In general, a wider range of an effective receptive field helps achieve better performance. Nonetheless, OctFormer achieves a chamfer distance and F-1 score of $7.45$ and $0.756$, which is worse than Octree U-Net by $1.31$ and $0.063$ respectively. This indicates that the OctFormer’s attention mechanism is less effective compared to an Octree U-Net architecture especially in the presence of latent 3D MAE, playing the similar role in the latent space.
@@
Table 7 shows the runtime performance of the baselines and our method. For a fair comparison, we run inference over the $50$ samples of the HOPE dataset and report the average time. For occupancy-based methods, we predict occupancy on object surfaces and occluded regions. Due to the memory-intensive nature of MCC [1]’s Transformer architecture, we run inference multiple times with the maximum chunk size of $10,000$ points. Our experiments demonstrate that implicit 3D representations used in POCO [1] and MCC [69] become slower when the voxel grid resolution is higher. Further, an autoregressive Transformer adopted in ShapeFormer [74] greatly increases the runtime. Conversely, the methods which leverage sparse voxel grids (e.g., MinkowskiNet [6], OCNN [66], and Ours) achieve much faster runtime thanks to efficient sparse 3D convolutions, and hierarchical pruning on predicted surfaces. Our method offers runtimes comparable to the fastest method, while implementing attention operations over the scene via latent 3D MAE, and achieving superior reconstruction.
@@
Figure 6: Qualitative comparison of OCNN [65] and our method. Our proposed latent 3D MAE helps predict globally consistent scene completion.